1. Find the Wronskian of the functions f(t)= (3e^t)sin t and g(t)= (e^t) cos t. Simplify your answer.

W(f,g)(t)= ________


2. Solve the initial value problem y′′+4y′+4y=0, y(−1)=4, y′(−1)=4.


1
Expert's answer
2020-07-09T19:50:12-0400


1)."W(f,g)(t)=\begin{vmatrix}
 3e^t\sin(t) & e^t\cos(t) \\
 (3e^t\sin(t))' & (e^t\cos(t))'
\end{vmatrix}=\begin{vmatrix}
 3e^t\sin(t) & e^t\cos(t) \\
 3\sqrt{2}e^t\sin(t+\pi/4) & \sqrt{2}e^t\cos(t+\pi/4)
\end{vmatrix}"

Thus,

"W(f,g)(t)=-3e^{2t}"


2).Given

"y\u2032\u2032+4y\u2032+4y=0, y(\u22121)=4, y\u2032(\u22121)=4."

Now, Auxiliary equation will be

"p^2+4p+4=0\implies (p+2)^2=0\implies p=-2"

Thus, general solution will be

"y=(a+bx)e^{-2x}"

Now,

"y'=-2(a+bx)e^{-2x}+be^{-2x}"

From, initial conditions we get,

"y(-1)=(a-b)e^2=4\\y'(-1)=-2(a-b)e^2+be^2=4\\
\implies b=12e^{-2},a=16e^{-2}"

Hence, the solution wil be

"y=(16+12x)e^{-2(x+1)}"